OpenStudy (anonymous):

PLEASE HELP: If a = 4, b = 16, c = 4.5, solve the triangle. (If there is no solution, give −1 as your answer.) a) Find C. (If there is more than one solution, give the larger one.) b) Find the corresponding angle A c) Find the corresponding angle B PLEASE HELP: If a = 4, b = 16, c = 4.5, solve the triangle. (If there is no solution, give −1 as your answer.) a) Find C. (If there is more than one solution, give the larger one.) b) Find the corresponding angle A c) Find the corresponding angle B @Mathematics

OpenStudy (mertsj):

cos B = (b^2-a^2-c^2)/-2ac cosB = (256-16-20.25)/-36 cosB = -6.1 Not possible. cosine cannot be smaller than -1 So the answer is -1